Raja received 21 NQOs (each option gives him the right to purchase 16 shares of stock for $24 per share) from his employer. At the time he started working, the stock price was $15 per share. Now that the share price is $24 per share, he intends to exercise all of the options using same-day sale. What are Raja's after tax proceeds from the sale if his marginal tax rate is 30 percent?
Answer #1

Raja's after tax proceeds from the sale is calculated below:

There are 21 Non-qualified stock options(NQOs) and option gives him the right to purchase 16 shares of stock

The Current share price at time of sale is $ 24 per share

Therefore, Sale Proceeds = (21*16)* $24 = $8,064

Cash required for exercising the options = Total Number of Shares* Strike Price

But Strike price is same with current stock price i.e.$24

So,Cash required for exercising the options = (21*16)* $24 = $8,064

Gain or loss on sale proceeds = $8,064 - $8,064 =0

No tax is charged on $0 income, so Raja's after tax proceeds from the sale is $0